OpenStudy (anonymous):

The distance between the upper and lower fixed points is 80 cm.Find the temperature on Celsius scale if the mercury level rises to a height 10.4 cm above the lower fixed point. What will be this temperature on the Fahrenheit scale?

OpenStudy (anonymous):

@RANE @rathanreddy @karatechopper

OpenStudy (anonymous):

sorry i am little weak in physics

OpenStudy (anonymous):

ohkk

OpenStudy (missmob):

yeah i didnt learn this yet cuz im new to this sub but ill try

OpenStudy (anonymous):

oh!! my bad luck :(

OpenStudy (missmob):

ill try though

OpenStudy (anonymous):

anyone?? help me

OpenStudy (missmob):

trying never hurts

OpenStudy (anonymous):

so lets try

OpenStudy (anonymous):

thank you :) @missMob i get it from my friend sits right side to me now... thank you for supporting me! @missMob and you my friend...lol

OpenStudy (missmob):

waait i know this onwee

OpenStudy (missmob):

just wait ill type

OpenStudy (missmob):

i was calculating

OpenStudy (anonymous):

okkk.. as your wish

OpenStudy (missmob):

Let us assume that the temperature is meausred betwene an upper point of 100C and a lower point of 0C. so, 80cm = 100C - 0C so, 1cm = 1.25C so, every cm represents 1.25C of temperature. thus, if the mercury rises to 10.4 cm, the temperature measured will be 10.4cm x 1.25C = 13C now, in Fahrenheit (F-32)/9 = C/5 so, F = 9C/5 + 32 = (9 x13) / 5 + 32 = 23.4 + 32 so, F = 55.4 F

OpenStudy (anonymous):

oh!!! just copied it from MERITNATION...

OpenStudy (anonymous):

forgot it.... but thanks for helping me

OpenStudy (missmob):

....................................................................................................................i calculated it too i just wanted to check the answer and i thpught that i have to write y not paste it LOL my bad soorry

OpenStudy (anonymous):

don't worry..... again thanks for helping me